Circular No. 8537

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 JY_{126} (CATALINA) Rik Hill reports his discovery of a comet on exposures taken on June

7.32 UT with the 0.68-m Schmidt telescope in the course of the Catalina Sky Survey, noting the object to be less distinct than surrounding stars and elongated along a northeast-southwest axis. C. W. Hergenrother, Lunar and Planetary Laboratory, writes that a co-added 1200-s R-band exposure with the University of Arizona 1.54-m Kuiper telescope shows the object to have a condensed, circular 15′′ coma and a thin, faint tail 35′′ long in p.a. 65^{o}. The observations were linked by the Minor Planet Center to the apparently asteroidal object 2005 JY_{126}, a Catalina discovery published on MPS 134992 (discovery observation given below).

2005 UT  α_{2000}  δ_{2000}  Mag.
May 12.31877  16^{h}13^{m}28.^{s}16  −4^{o}55′13′′.7  17.4

Additional astrometry (including prediscovery observations), the following orbital elements, and an ephemeris appear on MPEC 2005-L36.

T = 2006 Feb. 21.1195 TT  ω = 117.^{o}5648
e = 0.433597  Ω = 207.9705 }2000.0
q = 2.125789 AU  i = 20.2256
a = 3.753138 AU n^{o} = 0.1355540 P = 7.271 years

SUPERNOVA 2005cg E. S. Rykoff, University of Michigan, on behalf of the ROTSE collabo-

ration, reports the discovery of a supernova in unfiltered CCD images taken on June 2.04 (at mag ∼ 18.2) and 3.03 UT (mag ∼ 17.7) with the 0.45m ROTSE-IIIc telescope at the ‘High-Energy Stereoscopic Systems’ site in Namibia. SN 2005cg is located at α = 21^{h}10^{m}50^{s}.42, δ = +0^{o}12′07′′.4 (equinox 2000.0), which is 0′′.7 north and 0′′.4 west of the core of the apparent host galaxy (which the Sloan Digital Sky Survey gives as mag g′ = 19.7); nothing is visible at this location on ROTSE-IIIc images taken on May 7.14 (limiting mag ∼ 18.4). R. Quimby, University of Texas, adds that a spectrum (range 420–890 nm), obtained on June 3.40 with the 9.2-m Hobby-Eberly Telescope (+ Marcario Low-Resolution Spectrograph) under very poor conditions by M. Shetrone and S. Rostopchin, shows SN 2005cg to be a type-Ia supernova. Taking narrow emission lines at 677 and 501 nm to be Hα and Hβ from the host galaxy gives a redshift of 9290 km/s (yielding an absolute magnitude of −15.9 for the host galaxy).

2005 June 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green